OpenStudy (anonymous):

Find the slope and the y-intercept of the line. Write fractional answers in lowest terms 4y+9x+8=15+9x

jimthompson5910 (jim_thompson5910):

4y+9x+8=15+9x 4y=15+9x-9x-8 4y=7 y=7/4 Equation is now in the form y=mx+b where m=0 and b=7/4 (note: y=7/4 is the same as y=0x+7/4) So the slope is 0 and the y-intercept is (0, 7/4)